A 75 kg patient is prescribed rosuvastatin. 20 mg, once per day. Given the following PK parameters, predict the steady-state plasma concentration of the drug in this patient.
How long would it take to achieve steady-state?

Bioavailability = 20%
Vd = 134L
plasma protein binding = 88%
metabolism: CYP2C9; N-desmethyl rosuvastatin (major metabolite)
elimination: feces (90%)
t1/2 = 19 hrs

Please explain or show work if you can. Thanks!
Answer #1

Ans in the given problem the dose is low and dosing interval is also very long that is once per day so the steady state values will also be low.now the bioavailabilty is 20 % so 4 mg of drug is in systemic circulation.

Vd is 134L

Vd = amount of drug/concentration of drug in plasma

so conc of drug in plasma = amount of drug/ Vd

= 4/134= 0.03 mg/L

Steady state is achieved when rate of input of drug is equal to rate of elimination.Also steady state concentration is calculated as total exposure divided by time interval

0.03/24=0.0012 mg/Lhr

Time required to achieve the staedy state is 5 time the half life=5*19 hrs= 4 days.